# Quantum entanglement in phase space

### v2

@@ Introduction
Besides propagating optical and microwave fields where homodyne measurements are routinely implemented, in a wide range of other CV systems is lacking the possibility of performing quadrature measurements in a natural way. Examples include the motional states of trapped ions [3], the microwave cavity modes in cavity [4] and circuit quantum electrodynamics [5] and the vibration modes in circuit quantum acoustodynamics [6]. In these platforms, which are described by the Jaynes-Cummings model, a much more natural approach is to preform measurements of the Wigner function in phase space through displaced parity measurements [7].
It is thus natural to ask whether entanglement between two systems can be concluded directly from measurements of their joint Wigner function. Formally, it is known that for all separable states $\rho_{AB}=\sum_{\lambda}p(\lambda)\rho_{A}^{\lambda}\rho_{B}^{\lambda}$ the joint Wigner function factorizes as
@@
where $x^{\prime}=x^{\prime}(x,p),p^{\prime}=p^{\prime}(x,p)]$ is an arbitrary linear transformation of the coordinate system for $B$, see Eq. (8) and Fig. 1, and $R$ is an arbitrary region in phase space. We provide a method to derive the optimal transformation and show that known criteria [1, 2, 11, 12, 13] can be understood as specific choices. In particular, we prove that for Gaussian states criterion (I) is necessary and sufficient, as it coincides with Duan-Simon criteria, and that criterion (III) can be expressed as the measurement of the reduced Wigner function at a single point in the phase space of a transformed basis. We then show examples demonstrating how our criteria successfully detect entanglement in a variety of routinely prepared Gaussian and non-Gaussian CV states, such as two-mode squeezed states, Werner states and entangled cat states.
Interestingly, we prove that Criterion (I) has a close connection with the entanglement negativity and can thus be used as a bound on entanglement monotones. This originates from a connection between Criterion (I) and the Peres-Horodecki separability criterion [14, 15], which states that all separable states must have positive partial transpose (PPT). This follows from the fact that $\rho_{AB}^{T_{B}}=\sum_{\lambda}p(\lambda)\rho_{A}^{\lambda}(\rho_{B}^{\lambda})^{T}$ is a valid state with non-negative eigenvalues. A violation of this condition indicates entanglement. Such criterion is particularly powerful in CV systems, where bound entanglement is a rare phenomenon [16]. For entangled states that are not PPT, the entanglement negativity is a useful entanglement monotone defined via the negative eigenvalues of $\rho_{AB}^{T_{B}}$ [17, 18, 19, 20], specifically, $\mathcal{N}(\rho)=(\|\rho^{T_{B}}\|_{\text{tr}}-1)/2$. In the SM Sec. I [21], we show that the violation $v$ of criterion (I), quantified as the difference between both sides, provides a lower bound for entanglement negativity $\mathcal{N}(\rho)\geq\pi v$. For Schmidt-correlated states, both the PPT criterion [22] and criterion (I) are sufficient and necessary conditions. Moreover, the optimal $\pi v$ can even accurately provide the exact value of the entanglement negativity.
$\text{l.h.s}=\frac{1}{2\pi|\sin 2\theta|}\iint_{R}\left|\left\langle\frac{|\cos\theta|}{\sqrt{\pi}}D(x\cos\theta,p\cos\theta)\hat{\Pi}D^{\dagger}(x\cos\theta,p\cos\theta)\otimes\frac{|\sin\theta|}{\sqrt{\pi}}D(x^{\prime}\sin\theta,p^{\prime}\sin\theta)\hat{\Pi}D^{\dagger}(x^{\prime}\sin\theta,p^{\prime}\sin\theta)\right\rangle\right|\text{d}x\text{d}p,$ (3)
@@ Results
where $\bm{\Omega}=\bigoplus_{1}^{N}\bigl{(}\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\bigr{)}$ is the symplectic form and $\chi(\bm{a})=\operatorname{Tr}\left[\hat{\rho}\exp\left(i\hat{\bm{\xi}}^{\top}\bm{\Omega}\bm{\alpha}\right)\right]$ is the characteristic function. Crucially, the value of the Wigner function at one point in phase space can be directly measured experimentally by using the fact that it can be written as a displaced parity measurement [29, 30, 31], namely
@@
On the other hand, a determinant $\Delta=-1$ indicates a mirror reflection given by $\left(\begin{array}[]{cc}\cos(2\theta)&\sin(2\theta)\\ \sin(2\theta)&-\cos(2\theta)\end{array}\right)$. This reflection occurs along the axis which is rotated counterclockwise by an angle $\theta$ relative to the positive $x$-axis. Such transformation is associated to an anti-unitary operator, which leaves the absolute value of the inner product $|\langle x,y\rangle|$ invariant according to Wigner’s theorem [36]. As an example, time reversal $(x,p)\rightarrow(x,-p)$ is a typical anti-unitary operator.
@@
For example, taking the axis of the mirror reflection as the $x$-axis, such a transformation has already been used to generalize the Peres-Horodecki separability criterion [14, 15] to continuous variable systems [1], since the transpose operation of the density matrix $\hat{\rho}\rightarrow\hat{\rho}^{T}$ is equivalent to the mirror reflection $W(x,p)\rightarrow W(x,-p)$ in phase space. The Simon criterion detects entanglement by showing that the partially transposed state violates the uncertainty relation that separable states must satisfy [1]. For two-mode Gaussian states, when the Simon criterion is violated and the partially transposed state is sent through a beam splitter, one output mode exhibits an extremely localized Wigner function that would correspond to an unphysical purity exceeding unity, as detailed in SM Sec. III [21]. In contrast, our criterion (I) employs a fundamentally different approach: rather than testing uncertainty relation violations, it detects when the partially transposed state violates the upper bound of the Wigner function that all physical states must satisfy, in closer analogy to the original PPT criterion. This distinction, which we illustrate through a phase-space interpretation in SM Sec. III [21], explains why criterion (I) can outperform the Simon criterion for certain non-Gaussian states, as demonstrated in Fig. 8. Furthermore, for Gaussian states, the Simon criterion essentially utilizes the complete covariance matrix information, which fully determines the shape of the joint Wigner function (the first-order moments only affect the position of the Gaussian distribution). In contrast, criterion (I) relies solely on partial information from a specific phase-space slice, requiring less data than that needed for full reconstruction of the joint Wigner function, while still enabling reliable entanglement detection.
@@
Apart from mirror reflection along the $x$-axis, the criteria (I)-(III) we present are more general, since they consider any linear transformation $(x,p)\rightarrow(x^{\prime},p^{\prime})$ of the coordinate system that follows Eq. (8). In addition, our criteria do not require full information of the state. We give a proof of these criteria in the Methods section. We further show that criterion (I) is necessary and sufficient for Gaussian states in the Sec. V of SM [21].
@@
To some extent, we are using a single slice of the two-mode joint Wigner function to certify entanglement. It would be beneficial to perform an optimization to find the best slice. This approach would allow us to observe the maximum violation of our criteria. (i) To maximize criterion (I), we simultaneously optimize the transformation $(x,p)\rightarrow(x^{\prime},p^{\prime})$ and the parameter $\theta$ within the range $0<\theta<\pi$. The real parameters which determine the transformation Eq. (8) in Eq. (9), must satisfy the condition $\Delta=\pm 1$. (ii) To maximize criterion (II), we first optimize the transformation Eq. (9) and the parameter $\theta$ as previously described, across the entire plane to find the optimal correlation that maximally violates the criterion. Afterward, we focus on the region where the Wigner function is primarily distributed. In this step, we minimize the integral region $R$, while still ensuring that criterion (II) remains violated. This approach allows us to further reduce the measurements needed for criterion (II), which for $R=\mathbb{R}^{2}$ is equivalent to performing full tomography for a single mode, thereby saving measurement resources. (iii) To minimize criterion (III), we apply the same optimization of transformation $(x^{\prime},p^{\prime})$ as described previously.
@@
Testing criterion (III) can be performed in two ways. The first is to measure it in its original form, which is equivalent to performing single-mode tomography to obtain a full slice of the two-mode Wigner function. Alternatively, if the transformation $(x^{\prime},p^{\prime})$ corresponds to a unitary transformations acting on $\rho_{B}$, a second, potentially weaker method can be applied, requiring to measure only one point of the reduced Wigner function. Specifically, let us consider a beam-splitter transformation of the modes, $\hat{a}_{\pm}=(\hat{a}_{A}\pm\hat{a}_{B})/\sqrt{2}$. Using Eq. (15) with $\theta=\pi/4$ and $\left(x^{\prime}=\sqrt{2}x_{+}-x,p^{\prime}=\sqrt{2}p_{+}-p\right)$, we arrive at the special case $\iint_{-\infty}^{\infty}W_{AB}\left(x,p,x^{\prime},p^{\prime}\right)\text{d}x\text{d}p=\frac{1}{2}W_{+}\left(x_{+},p_{+}\right)$, which is half the Wigner function of one of the two modes resulting from the beam-splitter transformation. This means that, to violate criterion (III) it is enough to find a single point in the $(x_{+},p_{+})$ space where $W_{+}$ is negative [11, 12], see SM Sec. VI [21].
@@
Applications and comparison to known criteria.– In the following, we illustrate how criteria (I,II,III) are applied to several experimentally relevant scenarios. For a comparison between our entanglement criteria and nonlocality/steering criteria based on the Wigner function see SM Sec. VII [21].
@@
In the most general case of a TMST state, i.e. $1>\eta>0,r>0$, Duan-Simon and PPT criteria are still necessary and sufficient conditions for entanglement [1, 2, 14, 39]. Remarkably, we can prove that this is also true for our criterion (I), see SM Sec. V [21]. On the other hand, while our criterion (II) is still useful, the bound it provides is not tight. In fact, for Gaussian states criterion (II) is a specific case of criterion (I) with $\theta=\pi/4$ and $R=\mathbb{R}^{2}$, and therefore (I) is strictly stronger than (II). In Fig. 2, we illustrate this with a concrete example. The shaded region indicates TMST states that are entangled. These states are all detected by Simon’s criterion, which is violated above the black solid line, since it is necessary and sufficient for Gaussian states [1]. The same conclusion can be obtained from our criterion (I). In addition, our criterion (II) detects only a fraction of all TMST entangled states, as indicated by the regions above the dashed coloured lines. Note that these depend on the squeezing $s$, while the bound obtained from criterion (I) is independent of $s$. Contrary to (I,II), entanglement criteria in Refs. [11, 12] fail to detect entanglement for any TMST states, more precisely, any Gaussian state, since they only work for states with Wigner negativity.
@@
In summary, we have shown that our criteria based on the Wigner function provide bounds that are tight for the Werner states considered, since they coincides with the ones obtained from PPT.
Figure 3: Entanglement detection for dephased cat states Eq. (13). (a): Slice of the joint Wigner function defined by $W_{AB}(x,p,x,-p)$, for $\gamma=2$ and $\epsilon=1$. Yellow dashed circles indicate the smallest region $R$ needed to violate criterion (II). (b): Amount of violation for criterion (II) with $\theta=\pi/4$, i.e. $\iint_{-\infty}^{\infty}|W_{AB}(x,p,x,-p)|\text{d}x\text{d}p-1/(4\pi)$, and for PPT criterion, i.e. $|\lambda_{\min}(\rho^{T_{B}})|$, as a function of $\epsilon$ and for three values of $\gamma$. Entanglement can be revealed by both criteria whenever $\gamma,\epsilon>0$.
@@
where $|\psi\rangle=\mathcal{N}(|\gamma,\gamma\rangle+|-\gamma,-\gamma\rangle)$ is a two-mode superposition of coherent states $|\gamma\rangle$, i.e. a two-mode cat state, with $\mathcal{N}$ a normalization constant, and $\epsilon\in[0,1]$ determines the purity in terms of a dephasing of the superposition.
@@
To bridge the gap between our theoretical criteria and practical implementations with current technology, we provide a comprehensive error analysis from two perspectives in SM Sec. IX [21]. First, we examine the impact of grid spacing when using discrete-point measurements. When implementing our criteria experimentally, the continuous Wigner function must be sampled on a discrete grid through displaced parity measurements. We demonstrate that this discretization introduces minimal errors–even for coarse sampling with grid spacing $\delta=0.5$–confirming the robustness of our approach. Second, we present a measurement strategy specifically designed for our criteria. We introduce a measurement protocol that evaluates our criteria through scanning measurements uniformly distributed over the relevant phase-space region (for example, using Fermat spiral patterns [41] or randomized measurement). This approach is statistically equivalent to single-point displaced parity measurements, thereby significantly reducing data requirements. For the paradigmatic two-mode squeezed vacuum state, our method requires two orders of magnitude fewer measurements than the four-point Bell inequality test [8, 9] to achieve comparable confidence levels.
@@
Beyond the specific examples demonstrated, each criterion targets distinct classes of entangled states. Criteria (I) and (II) provide lower bounds on the entanglement negativity and the degree of CCNR criterion violation, respectively. Criterion (III), in the meanwhile, specifically detects entangled states that exhibit Wigner negativity.
@@
It would be interesting to investigate whether criteria similar to (I,II,III) can be found for discrete variable systems, using the Wigner function for spin states [42]. This would allow for alternative methods for the detection of entanglement between e.g. atomic ensembles [43, 44, 45, 46, 47]. We note that our approach readily extends beyond beamsplitters to encompass more general two-mode transformations of phase-space coordinates, including nonlinear operations such as two-mode squeezing.
Methods
Proof of criterion (I) – Consider an arbitrary transformation $(x^{\prime},p^{\prime})$ which can be determined by the real parameters $a,b,c,d,x_{0},p_{0}$. Based on these parameters, we then construct a new single-mode transformation $h:\left(\begin{array}[]{c}x\\ p\end{array}\right)\rightarrow\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\left(\begin{array}[]{c}x+\frac{X}{\cos\theta}\\ p+\frac{P}{\cos\theta}\end{array}\right)+\left(\begin{array}[]{c}x_{0}\sin\theta\\ p_{0}\sin\theta\end{array}\right)$ where $X,P$ are arbitrary constant coordinates. Then according to Eq. (9) we obtain
$\iint_{-\infty}^{\infty}\sum_{\lambda}p(\lambda)W_{A}^{\lambda}(x\cos\theta,p\cos\theta)$ (15)
$\times W_{h\left(B\right)}^{\lambda}\left(x\sin\theta-\frac{X}{\cos\theta},p\sin\theta-\frac{P}{\cos\theta}\right)\text{d}x\text{d}p$
$=$ $\iint_{-\infty}^{\infty}\sum_{\lambda}p(\lambda)W_{A}^{\lambda}\left(x\cos\theta+X\sin\theta,p\cos\theta+P\sin\theta\right)$
$\times W_{h\left(B\right)}^{\lambda}\left(x\sin\theta-X\cos\theta,p\sin\theta-P\cos\theta\right)\text{d}x\text{d}p\;,$
where in the last row we substituted $x\rightarrow x+X\tan\theta$ and similarly for $p$. This can be interpreted as follows: Perform a beam-splitter transformation $\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{array}\right)$ to the two modes $A$ and $h(B)$, and denote the resulting modes as $A^{\prime}:(x,p)$ and $B^{\prime}:(X,P)$. Notice that Eq. (15) is exactly the Wigner function of mode $B^{\prime}$, with mode $A^{\prime}$ integrated out. Therefore Eq. (15) should be upper-bounded by $1/(2\pi)$ [48]. Detailed proof can be found in Sec. IV of SM [21]. $\Box$
$\leqslant\sum_{\lambda}p(\lambda)\iint_{R}\left|W_{A}^{\lambda}(x\cos\theta,p\cos\theta)W_{f\left(\rho_{B}\right)}^{\lambda}(x\sin\theta,p\sin\theta)\right|\text{d}x\text{d}p$
$\leqslant\sum_{\lambda}p(\lambda)\sqrt{\left(\iint_{R}\left[W_{A}^{\lambda}(x\cos\theta,p\cos\theta)\right]^{2}\text{d}x\text{d}p\right)}$
$\quad\times\sqrt{\left(\iint_{R}\left[W_{f\left(\rho_{B}\right)}^{\lambda}(x\sin\theta,p\sin\theta)\right]^{2}\text{d}x\text{d}p\right)}$
$=\sum_{\lambda}p(\lambda)\sqrt{\left(\frac{\iint_{R}\left[W_{A}^{\lambda}(x,p)\right]^{2}\text{d}x\text{d}p}{\cos^{2}\theta}\right)\left(\frac{\iint_{R}\left[W_{f\left(\rho_{B}\right)}^{\lambda}(x,p)\right]^{2}\text{d}x\text{d}p}{\sin^{2}\theta}\right)}$
$\leqslant\frac{1}{4\pi|\cos\theta\sin\theta|}\;,$ (16)
where $R$ is an arbitrary region in phase space. Here, in the third-to-last row we used the Cauchy–Schwarz inequality, and in the last row we used Eq. (7). $\Box$
The physical motivation for criterion (III) coincides with fidelity calculations between two modes. For a pure product state, the fidelity between modes has a well-defined lower bound of zero. Since this expression is linear, it generalizes directly to mixed separable states. However, when the initial state is entangled, applying the same fidelity formula can yield negative values because the formula no longer represents a valid fidelity between modes for entangled states. This negative value indicates non-separability.
$\geqslant$ $0\;,$ (17)
where in the last row we used Eq.(6). $\Box$

### v3

@@ Introduction
Besides propagating optical and microwave fields where homodyne measurements are routinely implemented, in a wide range of other CV systems is lacking the possibility of performing quadrature measurements in a natural way. Examples include the motional states of trapped ions [3], the microwave cavity modes in cavity [4] and circuit quantum electrodynamics [5] and the vibration modes in circuit quantum acoustodynamics [6]. In these platforms, which are described by the Jaynes-Cummings model, a much more natural approach is to perform measurements of the Wigner function in phase space through displaced parity measurements [7].
It is thus natural to ask whether entanglement between two systems can be concluded directly from measurements of their joint Wigner function. Formally, it is known that for all separable states $\rho_{AB}=\sum_{\lambda}p(\lambda)\rho_{A}^{\lambda}\otimes\rho_{B}^{\lambda}$ the joint Wigner function factorizes as
@@
where $x^{\prime}=x^{\prime}(x,p),p^{\prime}=p^{\prime}(x,p)]$ is an arbitrary linear transformation of the coordinate system for $B$, see Eq. (7) and Fig. 1, and $R$ is an arbitrary region in phase space. The condition $\sin 2\theta\neq 0$ is assumed. We provide a method to derive the optimal transformation and show that known criteria [1, 2, 11, 12, 13] can be understood as specific choices. In particular, we prove that for Gaussian states criterion (I) is necessary and sufficient, as it coincides with Duan-Simon criteria, and that criterion (III) can be expressed as the measurement of the reduced Wigner function at a single point in the phase space of a transformed basis. We then show examples demonstrating how our criteria successfully detect entanglement in a variety of routinely prepared Gaussian and non-Gaussian CV states, such as two-mode squeezed states, Werner states and entangled cat states.
@@ Results
where $\bm{\Omega}=\bigoplus_{1}^{N}\bigl(\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\bigr)$ is the symplectic form and $\chi(\bm{\alpha})=\operatorname{Tr}\left[\hat{\rho}\exp\left(i\hat{\bm{\xi}}^{\top}\bm{\Omega}\bm{\alpha}\right)\right]$ is the characteristic function. Crucially, the value of the Wigner function at one point in phase space can be directly measured experimentally by using the fact that it can be written as a displaced parity measurement [14, 15, 16], namely
@@
On the other hand, a determinant $\Delta=-1$ indicates a mirror reflection given by $\left(\begin{smallmatrix}\cos(2\vartheta)&\sin(2\vartheta)\\ \sin(2\vartheta)&-\cos(2\vartheta)\end{smallmatrix}\right)$. This reflection occurs along the axis which is rotated counterclockwise by an angle $\vartheta$ relative to the positive $x$-axis. Such transformation is associated to an anti-unitary operator, which leaves the absolute value of the inner product $|\langle x,y\rangle|$ invariant according to Wigner’s theorem [21]. As an example, time reversal $(x,p)\rightarrow(x,-p)$ is a typical anti-unitary operator.
@@
For example, taking the axis of the mirror reflection as the $x$-axis, such a transformation has already been used to generalize the Peres-Horodecki separability criterion [23, 24] to continuous variable systems [1], since the transpose operation of the density matrix $\hat{\rho}\rightarrow\hat{\rho}^{T}$ is equivalent to the mirror reflection $W(x,p)\rightarrow W(x,-p)$ in phase space. The Simon criterion detects entanglement by showing that the partially transposed state violates the uncertainty relation that separable states must satisfy [1]. For two-mode Gaussian states, when the Simon criterion is violated and the partially transposed state is sent through a beam splitter, one output mode exhibits an extremely localized Wigner function that would correspond to an unphysical purity exceeding unity, as detailed in SM Sec. III [22]. In contrast, our criterion (I) employs a fundamentally different approach: rather than testing uncertainty relation violations, it detects when the partially transposed state violates the upper bound of the Wigner function that all physical states must satisfy, in closer analogy to the original PPT criterion. This distinction, which we illustrate through a phase-space interpretation in SM Sec. III [22], explains why criterion (I) can outperform the Simon criterion for certain non-Gaussian states (see Fig. 4 later in the text for a detailed comparison). Furthermore, for Gaussian states, the Simon criterion essentially utilizes the complete covariance matrix information, which fully determines the shape of the joint Wigner function (the first-order moments only affect the position of the Gaussian distribution). In contrast, criterion (I) relies solely on partial information from a specific phase-space slice, requiring less data than that needed for full reconstruction of the joint Wigner function, while still enabling reliable entanglement detection.
@@
Apart from mirror reflection along the $x$-axis, the criteria (I)-(III) we present are more general, since they consider any linear transformation $(x,p)\rightarrow(x^{\prime},p^{\prime})$ of the coordinate system that follows Eq. (7). In addition, our criteria do not require full information of the state. We give a proof of these criteria in the following. We further show that criterion (I) is necessary and sufficient for Gaussian states in the Sec. V of SM [22].
Proof of criterion (I) – Consider an arbitrary transformation $(x^{\prime},p^{\prime})$ which can be determined by the real parameters $a,b,c,d,x_{0},p_{0}$. Based on these parameters, we then construct a new single-mode transformation $h:(x,p)^{T}\rightarrow\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\left(x+\frac{X}{\cos\theta},p+\frac{P}{\cos\theta}\right)^{T}+\left(x_{0}\sin\theta,p_{0}\sin\theta\right)^{T}$ where $X,P$ are arbitrary constant coordinates. Then according to Eq. (8) we obtain
$\iint_{-\infty}^{\infty}\sum_{\lambda}p(\lambda)W_{A}^{\lambda}(x\cos\theta,p\cos\theta)W_{h\left(B\right)}^{\lambda}\left(x\sin\theta-\frac{X}{\cos\theta},p\sin\theta-\frac{P}{\cos\theta}\right)\text{d}x\text{d}p$ (11)
$=$ $\iint_{-\infty}^{\infty}\sum_{\lambda}p(\lambda)W_{A}^{\lambda}\left(x\cos\theta+X\sin\theta,p\cos\theta+P\sin\theta\right)W_{h\left(B\right)}^{\lambda}\left(x\sin\theta-X\cos\theta,p\sin\theta-P\cos\theta\right)\text{d}x\text{d}p\;,$
where in the last row we substituted $x\rightarrow x+X\tan\theta$ and similarly for $p$. This can be interpreted as follows: Perform a beam-splitter transformation $\left(\begin{smallmatrix}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{smallmatrix}\right)$ to the two modes $A$ and $h(B)$, and denote the resulting modes as $A^{\prime}:(x,p)$ and $B^{\prime}:(X,P)$. Notice that Eq. (11) is exactly the Wigner function of mode $B^{\prime}$, with mode $A^{\prime}$ integrated out. Therefore Eq. (11) should be upper-bounded by $1/(2\pi)$ [25]. Detailed proof can be found in Sec. IV of SM [22]. $\Box$
$\leqslant\sum_{\lambda}p(\lambda)\iint_{R}\left|W_{A}^{\lambda}(x\cos\theta,p\cos\theta)W_{f_{\theta}(\rho_{B}^{\lambda})}^{\lambda}(x\sin\theta,p\sin\theta)\right|\text{d}x\text{d}p$
$\leqslant\sum_{\lambda}p(\lambda)\sqrt{\left(\iint_{R}\left[W_{A}^{\lambda}(x\cos\theta,p\cos\theta)\right]^{2}\text{d}x\text{d}p\right)}\sqrt{\left(\iint_{R}\left[W_{f_{\theta}(\rho_{B}^{\lambda})}^{\lambda}(x\sin\theta,p\sin\theta)\right]^{2}\text{d}x\text{d}p\right)}$
$\leqslant\sum_{\lambda}p(\lambda)\sqrt{\left(\frac{\iint_{\mathbb{R}^{2}}\left[W_{A}^{\lambda}(x,p)\right]^{2}\text{d}x\text{d}p}{\cos^{2}\theta}\right)\left(\frac{\iint_{\mathbb{R}^{2}}\left[W_{f_{\theta}(\rho_{B}^{\lambda})}^{\lambda}(x,p)\right]^{2}\text{d}x\text{d}p}{\sin^{2}\theta}\right)}$
$\leqslant\frac{1}{4\pi|\cos\theta\sin\theta|}\;,$ (12)
where $R$ is an arbitrary region in phase space and we use $W_{B}^{\lambda}\left(x^{\prime}\sin\theta,p^{\prime}\sin\theta\right)=W_{f_{\theta}\left(\rho_{B}\right)}^{\lambda}(x\sin\theta,p\sin\theta)$. Here, in the third-to-last row we used the Cauchy–Schwarz inequality, and in the last row we used Eq. (6). $\Box$
The physical motivation for criterion (III) coincides with overlap calculations between two modes. For a pure product state, the fidelity between modes has a well-defined lower bound of zero. Since this expression is linear, it generalizes directly to mixed separable states. However, when the initial state is entangled, applying the same overlap formula can yield negative values because the formula no longer represents a valid fidelity between modes for entangled states. This negative value indicates non-separability.
$\geqslant$ $0\;,$ (13)
where in the last row we used Eq.(5). $\Box$
@@
To some extent, we are using a single slice of the two-mode joint Wigner function to certify entanglement. It would be beneficial to perform an optimization to find the best slice. This approach would allow us to observe the maximum violation of our criteria. (i) To maximize criterion (I), we simultaneously optimize the transformation $(x,p)\rightarrow(x^{\prime},p^{\prime})$ and the parameter $\theta$ within the range $0<\theta<\pi$ and $\theta\neq\pi/2$. The real parameters which determine the transformation Eq. (7) in Eq. (8), must satisfy the condition $\Delta=\pm 1$. (ii) To maximize criterion (II), we first optimize the transformation Eq. (8) and the parameter $\theta$ as previously described, across the entire plane to find the optimal correlation that maximally violates the criterion. Afterward, we focus on the region where the Wigner function is primarily distributed. In this step, we minimize the integral region $R$, while still ensuring that criterion (II) remains violated. This approach allows us to further reduce the measurements needed for criterion (II), which for $R=\mathbb{R}^{2}$ is equivalent to performing full tomography for a single mode, thereby saving measurement resources. (iii) To minimize criterion (III), we apply the same optimization of transformation $(x^{\prime},p^{\prime})$ as described previously.
@@
Testing criterion (III) can be performed in two ways. The first is to measure it in its original form, which is equivalent to performing single-mode tomography to obtain a full slice of the two-mode Wigner function. Alternatively, if the transformation $(x^{\prime},p^{\prime})$ corresponds to a unitary transformation acting on $\rho_{B}$, a second, potentially weaker method can be applied, requiring to measure only one point of the reduced Wigner function. Specifically, let us consider a beam-splitter transformation of the modes, $\hat{a}_{\pm}=(\hat{a}_{A}\pm\hat{a}_{B})/\sqrt{2}$. Using Eq. (11) with $\theta=\pi/4$ and $\left(x^{\prime}=\sqrt{2}x_{+}-x,p^{\prime}=\sqrt{2}p_{+}-p\right)$, we arrive at the special case $\iint_{-\infty}^{\infty}W_{AB}\left(x,p,x^{\prime},p^{\prime}\right)\text{d}x\text{d}p=\frac{1}{2}W_{+}\left(x_{+},p_{+}\right)$, which is half the Wigner function of one of the two modes resulting from the beam-splitter transformation. This means that, to violate criterion (III) it is enough to find a single point in the $(x_{+},p_{+})$ space where $W_{+}$ is negative [11, 12], see SM Sec. VI [22].
Comparison to known criteria and applications.– Interestingly, we prove that Criterion (I) has a close connection with the entanglement negativity and can thus be used as a bound on entanglement monotones. This originates from a connection between Criterion (I) and the Peres-Horodecki separability criterion [23, 24], which states that all separable states must have positive partial transpose (PPT). This follows from the fact that $\rho_{AB}^{T_{B}}=\sum_{\lambda}p(\lambda)\rho_{A}^{\lambda}\otimes(\rho_{B}^{\lambda})^{T}$ is a valid state with non-negative eigenvalues. A violation of this condition indicates entanglement. Such criterion is particularly powerful in CV systems, where bound entanglement is a rare phenomenon [26]. For entangled states that are not PPT, the entanglement negativity is a useful entanglement monotone defined via the negative eigenvalues of $\rho_{AB}^{T_{B}}$ [27, 28, 29, 30], specifically, $\mathcal{N}(\rho)=(\|\rho^{T_{B}}\|_{\text{tr}}-1)/2$. In the SM Sec. I [22], we show that the violation $v$ of criterion (I), quantified as the difference between both sides, provides a lower bound for entanglement negativity $\mathcal{N}(\rho)\geq\pi v$. For Schmidt-correlated states, both the PPT criterion [31] and criterion (I) are sufficient and necessary conditions. Moreover, the optimal $\pi v$ can even accurately provide the exact value of the entanglement negativity.
$\text{l.h.s}=\frac{1}{2\pi|\sin 2\theta|}$ $\iint_{R}\Bigl|\Bigl\langle\frac{|\cos\theta|}{\sqrt{\pi}}D(x\cos\theta,p\cos\theta)\hat{\Pi}D^{\dagger}(x\cos\theta,p\cos\theta)$
$\otimes\frac{|\sin\theta|}{\sqrt{\pi}}D(x^{\prime}\sin\theta,p^{\prime}\sin\theta)\hat{\Pi}D^{\dagger}(x^{\prime}\sin\theta,p^{\prime}\sin\theta)\Bigr\rangle\Bigr|\text{d}x\text{d}p,$
@@
In the following, we illustrate how criteria (I,II,III) are applied to several experimentally relevant scenarios. For a comparison between our entanglement criteria and nonlocality/steering criteria based on the Wigner function see SM Sec. VII [22].
@@
In the most general case of a TMST state, i.e. $1>\eta>0,r>0$, Duan-Simon and PPT criteria are still necessary and sufficient conditions for entanglement [1, 2, 23, 40]. Remarkably, we can prove that this is also true for our criterion (I), see SM Sec. V [22]. On the other hand, while our criterion (II) is still useful, the bound it provides is not tight. In fact, for Gaussian states criterion (II) is a specific case of criterion (I) with $\theta=\pi/4$ and $R=\mathbb{R}^{2}$, and therefore (I) is strictly stronger than (II). In Fig. 2, we illustrate this with a concrete example. The shaded region indicates TMST states that are entangled. These states are all detected by Simon’s criterion, which is violated above the black solid line, since it is necessary and sufficient for Gaussian states [1]. The same conclusion can be obtained from our criterion (I). In addition, our criterion (II) detects only a fraction of all TMST entangled states, as indicated by the regions above the dashed coloured lines. Note that these depend on the squeezing $s$, while the bound obtained from criterion (I) is independent of $s$. In contrast to (I,II), entanglement criteria in Refs. [11, 12] fail to detect entanglement for any TMST states, more precisely, any Gaussian state, since they only work for states with Wigner negativity.
@@
In summary, we have shown that our criteria based on the Wigner function provide bounds that are tight for the Werner states considered, since they coincide with the ones obtained from PPT.
Figure 3: Entanglement detection for dephased cat states Eq. (16). (a) Slice of the joint Wigner function defined by $W_{AB}(x,p,x,-p)$, for $\gamma=2$ and $\epsilon=1$. Yellow dashed circles indicate the smallest region $R$ needed to violate criterion (II). (b) Amount of violation for criterion (II) with $\theta=\pi/4$, i.e. $\iint_{-\infty}^{\infty}|W_{AB}(x,p,x,-p)|\text{d}x\text{d}p-1/(4\pi)$, and for PPT criterion, i.e. $|\lambda_{\min}(\rho^{T_{B}})|$, as a function of $\epsilon$ and for three values of $\gamma$. Entanglement can be revealed by both criteria whenever $\gamma,\epsilon>0$.
@@
where $|\psi\rangle=\mathcal{N}(|\gamma,\gamma\rangle+|-\gamma,-\gamma\rangle)$ is a two-mode superposition of coherent states $|\gamma\rangle$, i.e. a two-mode cat state, with $\mathcal{N}$ a normalization constant, and $\epsilon\in[0,1]$ determines the mixing weight in terms of a dephasing of the superposition.
@@
To bridge the gap between our theoretical criteria and practical implementations with current technology, we provide a comprehensive error analysis from two perspectives in SM Sec. IX [22]. First, we examine the impact of grid spacing when using discrete-point measurements. When implementing our criteria experimentally, the continuous Wigner function must be sampled on a discrete grid through displaced parity measurements. We demonstrate that this discretization introduces minimal errors–even for coarse sampling with grid spacing $\delta=0.5$–confirming the robustness of our approach. Second, we present a measurement strategy specifically designed for our criteria. We introduce a measurement protocol that evaluates our criteria through scanning measurements uniformly distributed over the relevant phase-space region (for example, using Fermat spiral patterns [43] or randomized measurement). This approach is statistically equivalent to single-point displaced parity measurements, thereby significantly reducing data requirements. For the paradigmatic two-mode squeezed vacuum state, our method requires two orders of magnitude fewer measurements than the four-point Bell inequality test [8, 9] to achieve comparable confidence levels. We further illustrate in SM Sec. IX [22] that, when the total number of measurements is expressed as $N_{\mathrm{tot}}=M\cdot N_{\mathrm{shot}}$, where $M$ denotes the number of measurement rounds and $N_{\mathrm{shot}}$ the number of shots per round, the randomized protocol requires only $M=1$ for criteria (I) and (III), and $M=2$ for criterion (II). This represents a significant reduction compared to full two-mode tomography, where $M$ typically scales as $D^{4}$ with $D$ being the effective truncation dimension of each mode.
@@
Beyond the specific examples demonstrated, each criterion targets distinct classes of entangled states. Criteria (I) and (II) provide lower bounds on the entanglement negativity and the degree of CCNR criterion violation, respectively. Criterion (III), in the meanwhile, specifically detects entangled states that exhibit Wigner negativity. For a detailed comparison, see Fig. 4.
@@
It would be interesting to investigate whether criteria similar to (I,II,III) can be found for discrete variable systems, using the Wigner function for spin states [44]. This would allow for alternative methods for the detection of entanglement between e.g. atomic ensembles [45, 46, 47, 48, 49]. We note that our approach readily extends beyond beamsplitters to encompass more general two-mode transformations of phase-space coordinates, including operations such as two-mode squeezing.